# Cross-Resolution SAR Target Detection Using Structural Hierarchy Adaptation and Reliable Adjacency Alignment

### v1

@@ IV-B Structure-induced Hierarchical Feature Adaptation
[…]
@@ IV-C Reliable Structural Adjacency Alignment
[…]
@@ IV-C2 Secure Adjacency Factor $\hat{\gamma}$
It is used to select the secure adjacency set of each target domain instance $x^{T}$, making it suitable for adjacency semantic alignment. We hypothesize that both structural and perceptual consistency are essential for a secure adjacency set. 1) Structural consistency: instances within the secure adjacency set exhibit structural consistency. 2) Perceptual consistency: instances within the secure adjacency set have similar levels of reliability. If both conditions are satisfied, we consider the adjacency set as a secure set, indicating the adjacency set has high-level structural and perceptual consistency. Based on these two hypotheses, the RSAA $\hat{\gamma}\left({x_{i}^{T}}\right)$ is defined from both structural and perceptual perspectives:
@@
${\tilde{D}_{ST}}=\left\{{{d_{ST}}\left({i,j}\right)|i,j\in\left|{{R_{r}}\left({x_{i}^{T}}\right)}\right|}\right\}$ is the set of scattering structure distances within ${R_{r}}\left({x_{i}^{T}}\right)$. The uncertainty $u(x_{i})$ serves as a measure of the prediction reliability for neighbors within the adjacency set. ${\left\|\cdot\right\|_{1}}$ is Manhattan distance. $\epsilon$ is a small constant to avoid division by zero.
@@ IV-C3 Secure Adjacency Set Division
The smaller ${{\hat{\gamma}}_{ST}}\left({x_{i}^{T}}\right)$ and ${{\hat{\gamma}}_{u}}\left({x_{i}^{T}}\right)$ indicate higher structural and semantic consistency within the $r$-adjacency set $R_{r}(x_{i}^{T})$. Accordingly, the adjacency set is divided as the secure adjacency set $R_{r}^{se}$. Conversely, large ${{\hat{\gamma}}_{ST}}\left({x_{i}^{T}}\right)$ and ${{\hat{\gamma}}_{u}}\left({x_{i}^{T}}\right)$ values suggest a risky adjacency set, which is discarded from semantic alignment. To avoid relying on fixed thresholds, we introduce an adaptive strategy for secure adjacency set selection:
$R_{r}^{ST}$ =$=\left\{{{R_{r}}\left({x_{i}^{T}}\right)|{{\hat{\gamma}}_{ST}}\left({x_{i}^{T}}\right)\leq{\mu_{ST}}+{\lambda_{se}}\cdot{\sigma_{ST}}}\right\},$ (17)
$R_{r}^{u}$ =$=\left\{{{R_{r}}\left({x_{i}^{T}}\right)|{{\hat{\gamma}}_{u}}\left({x_{i}^{T}}\right)\leq{\mu_{u}}+{\lambda_{se}}\cdot{\sigma_{u}}}\right\}.$
$\mu$ and $\sigma$ represent the mean and standard deviation of $\hat{\gamma}$, respectively. ${\lambda}_{\textit{se}}$ is a scaling hyperparameter. Finally, the secure adjacency set $R_{r}^{se}$ can be obtained as the intersection of the structural and perceptual constraints:
@@ IV-C4 Soft Semantic Alignment
To enhance the discriminability of detector in the target domain, the secure adjacency set $R_{r}^{se}$ is further partitioned as secure foreground adjacency set $R_{r}^{fg}$ and secure background adjacency set $R_{r}^{bg}$ according to the probabilities of source adjacency instances:
$R_{r}^{fg}$ =$=\left\{{R_{r}^{se}\left({x_{i}^{T}}\right)|\mathop{\min}\limits_{x_{j}^{S}\in{N_{r}}\left({x_{i}^{T}}\right)}p_{j}^{S}\geq 0.5}\right\},$ (19)
$R_{r}^{bg}$ =$=\left\{{R_{r}^{se}\left({x_{i}^{T}}\right)|\mathop{\min}\limits_{x_{j}^{S}\in{N_{r}}\left({x_{i}^{T}}\right)}p_{j}^{S}<0.5}\right\}.$
@@ IV-D Overall Training Objective
In summary, to achieve cross-resolution SAR target detection, the overall optimization objective of our method can be expressed as:
@@ V-B Implementation Details
All experiments were conducted on a server equipped with an NVIDIA A6000 GPU with 48GB of memory. During the training phase, labeled source domain data and unlabeled target domain data were jointly input into the target detection model for training, and all methods were trained for 10 epochs. The initial learning rate was set as $0.002$, and decreased to $0.001$ after 5 epochs. In the testing phase, the target domain data were used to evaluate the detection performance. The evaluation metrics we adopted include VOC mAP, F1 score, Precision, and Recall. In all experiments, the IoU threshold for calculating mAP was set to 0.5. Hyper-parameter ${\lambda}_{\textit{SHFA}}=0.5$, ${\lambda}_{\textit{RSAA}}=0.5$, ${\lambda}_{\textit{se}}=-1$, $k=30$ $n_{a}=5$, $r=5$ for Aircraft tasks, and $r=3$ for Vehicle tasks. Details are discussed in Section V-F.
@@ V-C Comparisons and Analysis
To ensure the fairness of the experiments, all comparison methods were implemented using the Faster RCNN detector, and comparative experiments were conducted based on the four cross-resolution target detection tasks shown in Table II.
@@ V-C1 Quantitative comparisons
From the quantitative results in Table III and Table IV, it is clear that the performance of Faster RCNN without domain adaptation is significantly lower than that of other methods. This demonstrates that discrepancies across resolutions can lead to a marked decline in SAR target detection performance. Moreover, five cross-domain detection methods perform significantly better than the baseline Faster RCNN, validating the feasibility and effectiveness of domain adaptation-based approaches in cross-domain target detection tasks.
According to Table III, HSANet and DAF perform poorly in the Aircraft LR→HR and Aircraft HR→LR tasks. This is because these cross-resolution detection tasks involve a greater diversity in target scales, which increases the challenge of cross-resolution target detection. In contrast, the proposed CR-Net outperforms the other four SOTA cross-domain detection methods across all evaluation metrics. This not only highlights the effectiveness of CR-Net in cross-resolution detection tasks, but also further validates its robustness and performance advantages in addressing resolution distribution discrepancies. In the Aircraft LR→HR task, CR-Net achieves over a 25% improvement in Recall, Precision, and F1 score compared to the second-best method, GPANet. For the Aircraft LR→HR task, our method achieves over a 28% improvement in Precision, while for the Aircraft HR→LR task, the improvement is approximately 13%. These results demonstrate that our method significantly enhances the detector’s discriminability. This is achieved by aligning feature distributions effectively by SHFA and effectively selecting high-quality discriminative instances by RSAA.
As for Vehicle LR→HR and Vehicle HR→LR results from Table IV, the mAP scores are similar to F1 scores. This similarity can be attributed to the relatively uniform scale distribution of vehicles in the dataset, which reduces the challenge of bounding box regression. Nevertheless, our method still achieves the best performance compared to other methods in both tasks. Specifically, in the Vehicle LR→HR task, our method achieves 0.713 F1 score, outperforming the second-best method, IDA, by approximately 12%. This substantial improvement demonstrates that our approach is not only suitable for addressing resolution discrepancies in meter-level datasets, but also effective for cross-resolution target detection in sub-meter, very high-resolution datasets.
TABLE III: Statistic comparisons with SOTA UDA methods on aircraft Tasks. The bold figure denotes the best score. The underline indicates the second-best performance.
@@
TABLE IV: Statistic comparisons with SOTA UDA methods on vehicle Tasks. The bold figure denotes the best score. The underline indicates the second-best performance.
@@ V-C2 Qualitative comparisons
The Precision-Recall (PR) curves of five methods across four different cross-domain detection tasks are shown in Fig. 8. The PR curves show that the area under the curve (AUC) of our method is significantly larger across all tasks, demonstrating its superiority. Additionally, the PR curve of the proposed method is more stable compared to other methods, indicating that the detection precision remains consistent under varying recall rates, highlighting stronger generalization capabilities. Moreover, to provide an intuitive comparison, the visualization results of detection are shown in Fig 10, Fig 12, Fig 15, and Fig 18. From these visualizations, although the other four methods can detect most targets, they also exhibit a large number of false positives and missed detections. In contrast, the proposed method significantly reduces false positives and missed detections. This improvement is particularly evident in complex scenarios, where the proposed method demonstrates stronger robustness and adaptability. From the aircraft detection results shown in Fig. 10, our method demonstrates remarkable discriminative performance even in complex contextual backgrounds, such as near jet bridges. Compared with other four methods that rely solely on adversarial feature alignment, our method enhances the discriminability by leveraging reliable source discrimination knowledge on top of structure-aware feature adaptation. This improvement enables our method to achieve superior performance in complex background scenarios.
@@ V-E2 Effective Analysis of RSAA Module
According to the results shown in Fig. 11 (a), when the RSAA module is not used and features are directly aligned using L1 loss, the F1 score drops sharply after 1000 training iterations. When the reliable instance factor $\hat{\eta}$ is removed (setting $k=0$), the F1 score initially increases but then decreases as training iterations progress. In contrast, when larger values ($k=30$ or $k=40$) are set, the F1 score steadily increases during training and shows significant improvement compared to the case without $\hat{\eta}$. This indicates that the reliable instance factor plays a positive role in mitigating error accumulation during the adjacency alignment process. Fig. 11 (b) illustrates the impact of the secure adjacency factor for iterative adjacency semantic alignment process. According to the definition of $\hat{\gamma}$ in Eq. 15, the smaller ${\lambda}_{\textit{se}}$ indicates the stronger the filtering effect on the risky set. Based on the results in Fig. 11, when $\hat{\gamma}$ is removed, the model experiences a sharp collapse in the F1 score after approximately 2000 iterations during the adjacency semantic alignment process. In contrast, when $\hat{\gamma}$ is introduced, the F1 score shows a gradual upward trend throughout the iterations. This demonstrates that the secure adjacency factor $\hat{\gamma}$ helps to filter out unreliable risky sets, preventing the propagation of risky information, which is critical for reliable adjacency alignment. To acquire satisfactory performance, the hyper-parameters are set to $k=30$ and $\lambda_{\text{Secmask}}=-1$ by default, as indicated by the results in Fig. 11.
@@ V-F1 Discussions on Hyperparameter of SHFA
Fig. 14 illustrates the effect of the number of typical structures $n_{a}$ (defined in Section IV-B), with ${\lambda}_{\textit{SHFA}}=0.5$. For both the Aircraft LR→HR and Vehicle LR→HR tasks, the results demonstrate that the F1 score and mAP vary with $n_{a}$. When $n_{a}$ is set to 0, the SHFA module degenerates into a conventional single-adversarial domain adaptation model, resulting in the lowest F1 and mAP values. As $n_{a}$ increases, more hierarchical discriminators are introduced into SHFA module, leading to improvements in both F1 and mAP. This indicates that assigning features to corresponding discriminators based on structural similarity can enhance performance. The best results are achieved when $n_{a}=5$, where both F1 and mAP reach their peaks. Further increasing $n_{a}$ does not yield additional improvements and may even cause a slight decrease in performance. Therefore, it is recommended to set $n_{a}$ to around 5 to ensure optimal performance.

### v2

@@ IV-B Structure-induced Hierarchical Feature Adaptation
Specifically, based on the scattering structure distance $d_{ST}$ introduced earlier, we use clustering methods to obtain $n_{a}$ structure anchors $\{A_{j}\}_{j=1}^{n_{a}}$, representing typical structures observed in the source domain data. Examples of these typical structures are shown in the Fig. 7. These typical structures reflect the main modes of the source domain feature distribution (visualized in Fig 6) and provide structural guidance for adapting domain features. Given the scattering point set $P_{i}$ of the $i\text{-th}$ instance, we can calculate the structural similarity $sim_{i,j}$ between this instance and each structure anchor $A_{j}$, which quantifies the relationship between the $i\text{-th}$ instance and the $j\text{-th}$ typical structures:
@@ IV-C Reliable Structural Adjacency Alignment
Given a target domain instance $x^{T}\in\mathcal{X}^{T}$, we define the $r\text{-adjacency set}$ of $x^{T}$ as ${R_{r}}\left({{x^{T}}}\right)=\left\{{{x^{T}}}\right\}\cup\left\{{{N_{r}}\left({{x^{T}}}\right)}\right\}$, where $N_{r}(x^{T})$ is the $r$ nearest source domain neighbors of the target domain instance $x^{T}$. Specifically:
@@ IV-C2 Secure Adjacency Factor $\hat{\gamma}$
It is used to select the secure adjacency set of each target domain instance $x^{T}$, making it suitable for adjacency semantic alignment. We hypothesize that both structural and perceptual consistency are essential for a secure adjacency set. 1) Structural consistency: instances within the secure adjacency set exhibit structural consistency. 2) Perceptual consistency: instances within the secure adjacency set have similar levels of reliability. If both conditions are satisfied, we consider the adjacency set as a secure set, indicating the adjacency set has high-level structural and perceptual consistency. Based on these two hypotheses, the secure adjacency factor $\hat{\gamma}\left({x_{i}^{T}}\right)$ is defined from both structural and perceptual perspectives:
@@
Here, ${{\hat{\gamma}}_{ST}}\left({x_{i}^{T}}\right)$ and ${{\hat{\gamma}}_{u}}\left({x_{i}^{T}}\right)$ respectively characterize structural consistency and perceptual consistency for the $r\text{-adjacency set}$ of $x^{T}_{i}$. ${\tilde{D}_{ST}}=\left\{d_{ST}\left(i,j\right)\mid i,j\in|{R_{r}}\left(x_{i}^{T}\right)|\right\}$ denotes the set of scattering structure distances between all pairs of elements within the $r$-adjacency set ${R_{r}}\left(x_{i}^{T}\right)$. The uncertainty $u(x_{i})$ quantifies the prediction reliability for the neighbors in the adjacency set. $\left\|\cdot\right\|_{1}$} is Manhattan distance. $\epsilon$ is a small constant to avoid division by zero.
@@ IV-C3 Secure Adjacency Set Division
Smaller values of ${{\hat{\gamma}}_{ST}}\left(x_{i}^{T}\right)$ and ${{\hat{\gamma}}_{u}}\left(x_{i}^{T}\right)$ indicate higher structural and perceptual consistency within the $r$-adjacency set $R_{r}(x_{i}^{T})$. Under such conditions, the adjacency set is categorized as a secure adjacency set $R_{r}^{se}$, which is used for reliable semantic alignment. In contrast, larger values of ${{\hat{\gamma}}_{ST}}\left(x_{i}^{T}\right)$ and ${{\hat{\gamma}}_{u}}\left(x_{i}^{T}\right)$ imply low consistency, and such risky adjacency sets are excluded from semantic alignment.
Therefore, to ensure that the selected adjacency set satisfies both structural and perceptual consistency, we define secure adjacency set $R_{r}^{se}$ as the intersection of the structurally consistent set $R_{r}^{ST}$ and the perceptually consistent set $R_{r}^{u}$:
@@
Here, $R_{r}^{ST}$ and $R_{r}^{u}$ are selected from $R_{r}$ by applying the structural consistency constraint ${{\hat{\gamma}}_{ST}}\left(x_{i}^{T}\right)$ and the perceptual consistency constraint ${{\hat{\gamma}}_{u}}\left(x_{i}^{T}\right)$, respectively. To enhance the flexibility of the selection process, we adopt an adaptive selection strategy:
$R_{r}^{ST}$ $=\left\{{{R_{r}}\left({x_{i}^{T}}\right)|{{\hat{\gamma}}_{ST}}\left({x_{i}^{T}}\right)\leq{\mu_{ST}}+{\lambda_{se}}\cdot{\sigma_{ST}}}\right\},$ (18)
$R_{r}^{u}$ $=\left\{{{R_{r}}\left({x_{i}^{T}}\right)|{{\hat{\gamma}}_{u}}\left({x_{i}^{T}}\right)\leq{\mu_{u}}+{\lambda_{se}}\cdot{\sigma_{u}}}\right\}.$
$\mu_{ST}$ and $\sigma_{ST}$ are the mean and standard deviation of $\hat{\gamma}_{ST}$ over the training batch. Similarly, $\mu_{u}$ and $\sigma_{u}$ are defined for $\hat{\gamma}_{u}$. ${\lambda}_{\textit{se}}$ is a scaling hyperparameter. Following this strategy, all $r$-adjacency set $R_{r}(x_{i}^{T})$ of $x_{i}^{T}\in\mathcal{X}^{T}$ that meet the meet the constraints in Eq.18 are included in the secure adjacency set $R_{r}^{se}$.
@@ IV-C4 Soft Semantic Alignment
To enhance the detector’s discriminability in the target domain, we use the reliable instance factor $\hat{\eta}$ and the secure adjacency set $R_{r}^{se}$ to guide soft semantic alignment, thereby transferring semantic information from the source domain to the target domain. Furthermore, we encourage the separation of feature distributions between foreground and background instances to further improve the model’s discriminability. Specifically, the secure adjacency set $R_{r}^{se}$ is partitioned as secure foreground adjacency set $R_{r}^{fg}$ and secure background adjacency set $R_{r}^{bg}$ according to the probabilities $p_{j}$ of neighbor instances $x_{j}\in N_{r}(x_{i}^{T})$:
$R_{r}^{fg}$ $=\left\{{R_{r}^{se}\left({x_{i}^{T}}\right)|\mathop{\min}\limits_{x_{j}\in{N_{r}}\left({x_{i}^{T}}\right)}p_{j}\geq 0.5}\right\},$ (19)
$R_{r}^{bg}$ $=\left\{{R_{r}^{se}\left({x_{i}^{T}}\right)|\mathop{\min}\limits_{x_{j}\in{N_{r}}\left({x_{i}^{T}}\right)}p_{j}<0.5}\right\}.$
@@ IV-D Overall Training Objective
In summary, to achieve cross-resolution SAR target detection, the overall optimization objective of CR-Net is expressed as:
@@ V-B Implementation Details
All experiments were conducted on a server equipped with an NVIDIA A6000 GPU with 48GB of memory. During the training phase, labeled source domain data and unlabeled target domain data were jointly input into the target detection model for training 10 epochs. The initial learning rate was set as $0.002$, and decreased to $0.001$ after 5 epochs. In the testing phase, the target domain data were used to evaluate the detection performance. The evaluation metrics we adopted include VOC mAP, F1 score, Precision, and Recall. In all experiments, the IoU threshold for calculating mAP was set to 0.5. Hyper-parameter ${\lambda}_{\textit{SHFA}}=0.5$, ${\lambda}_{\textit{RSAA}}=0.5$, ${\lambda}_{\textit{se}}=-1$, $k=30$ $n_{a}=5$, $r=5$ for Aircraft tasks, and $r=3$ for Vehicle tasks. Details are discussed in Section V-F.
@@ V-C Comparisons and Analysis
In the comparative experiments, we compared seven SOTA domain adaptation methods, including Domain Adaptive Faster RCNN [68] (DAF), Graph-induced Prototype Alignment Network [71] (GPANet), Imbalanced Discriminant Alignment [7] (IDA), Hierarchical Similarity Alignment Neural Network [8] (HSANet), Foreground Instance Enhancement Network [72] (FIENet), Class-Aware Teacher [52] (CAT), and Masked Image Consistency [73] (MIC). Among these methods, only FIENet is based on the FCOS [74] detection model, while all other methods use the Faster RCNN detector with ResNet-101. To ensure fair comparison, all experiments were conducted with the same input image size and identical data augmentation strategies. Comparative experiments were carried out on the four cross-resolution target detection tasks presented in Table II
@@ V-C1 Quantitative comparisons
From the quantitative results in Table III and Table IV, it is clear that the performance of Faster RCNN without domain adaptation is significantly lower than that of other methods. This demonstrates that discrepancies across resolutions can lead to a substantial decline in SAR target detection performance. Moreover, domain adaptive detection methods consistently outperform the baseline Faster RCNN, thus validating the feasibility and effectiveness of domain adaptation-based approaches for cross-domain target detection tasks.
According to the meter-level cross-resolution results in Table III, HSANet and DAF perform poorly in the Aircraft LR→HR and Aircraft HR→LR tasks. This is because these cross-resolution detection tasks involve a greater diversity in target scales, which increases the challenge of cross-resolution target detection. In contrast, the proposed CR-Net outperforms the other detection methods across all evaluation metrics. This not only highlights the effectiveness of CR-Net in cross-resolution detection, but also further demonstrates its robustness and performance advantages in addressing resolution distribution discrepancies. For the meter-level cross-resolution Aircraft LR→HR and HR→LR tasks, CR-Net achieves improvements of more than 14 % in Recall, mAP, and F1 score, and approximately 4 % in Precision, compared with the second-best method. These results indicate that our method significantly enhances the detector’s transferability and discriminability for the cross-resolution detection task. This is achieved by aligning feature distributions effectively by SHFA and transferring high-quality, discriminative semantics via RSAA.
As for the sub-meter level cross-resolution results (Vehicle LR→HR and Vehicle HR→LR) in Table IV, the mAP scores are closer to the F1 scores. This can be attributed to the relatively uniform scale distribution of vehicles in the dataset, which reduces the challenge of bounding box regression during the detection process. Nevertheless, our method still achieves the best performance in Recall, mAP, and F1 in both tasks. Specifically, in the Vehicle HR→LR task, our method achieves 0.652 mAP and 0.701 F1 score, outperforming the second-best method, MIC, by approximately 16 %. Although, in the Vehicle LR→HR task, our method’s Precision is slightly lower than IDA by 1.2 %, it surpasses IDA by at least 10 % in Recall, F1, and mAP metrics, indicating that our method achieves a better balance between recall and precision.
Based on the above results and analysis, our method is not only suitable for addressing resolution discrepancies in meter-level datasets, but is also effective for cross-resolution target detection in sub-meter, very high-resolution datasets.
TABLE III: Statistic comparisons with SOTA UDA methods on aircraft Tasks. The bold figure denotes the best score. The underline indicates the second-best performance. The symbol ‘*’ denotes the detector based on the FCOS model.
@@
TABLE IV: Statistic comparisons with SOTA UDA methods on vehicle Tasks. The bold figure denotes the best score. The underline indicates the second-best performance. The symbol ‘*’ denotes the detector based on the FCOS model.
@@ V-C2 Qualitative comparisons
The Precision-Recall (PR) curves of comparison methods across different cross-domain detection tasks are shown in Fig. 9. The PR curves show that the area under the curve (AUC) of our method is significantly larger across all tasks, demonstrating its superiority. Additionally, the PR curve of the proposed method is more stable compared to other methods, indicating that the detection precision remains consistent under varying recall rates, highlighting stronger generalization capabilities.
Moreover, to provide an intuitive comparison, the visualization results of detection are shown in Fig 11, Fig 13, Fig 16, and Fig 19. From these visualizations, although the other comparison methods can detect most targets, they also exhibit a large number of false positives and missed detections. In contrast, the proposed method significantly reduces false positives and missed detections. This improvement is particularly evident in complex scenarios, where the proposed method demonstrates stronger robustness and adaptability. From the aircraft detection results shown in Fig. 11, our method demonstrates remarkable discriminative performance even in complex contextual backgrounds, such as near jet bridges. Compared with other methods that rely solely on adversarial feature alignment, our method enhances the discriminability by leveraging reliable source discrimination knowledge transfer through RSAA. This improvement enables our method to achieve superior performance in complex background scenarios.
@@ V-E2 Effective Analysis of RSAA Module
According to the results shown in Fig. 12 (a), when the RSAA module is not used and features are directly aligned using L1 loss, the F1 score drops sharply after 1000 training iterations. When the reliable instance factor $\hat{\eta}$ is removed (setting $k=0$), the F1 score initially increases but then decreases as training iterations progress. In contrast, when larger values ($k=30$ or $k=40$) are set, the F1 score steadily increases during training and shows significant improvement compared to the case without $\hat{\eta}$. This indicates that the reliable instance factor plays a positive role in mitigating error accumulation during the adjacency alignment process. Fig. 12 (b) illustrates the impact of the secure adjacency factor for iterative adjacency semantic alignment process. According to the definition of $\hat{\gamma}$ in Eq. 16, the smaller ${\lambda}_{\textit{se}}$ indicates the stronger the filtering effect on the risky set. Based on the results in Fig. 12, when $\hat{\gamma}$ is removed, the model experiences a sharp collapse in the F1 score after approximately 2000 iterations during the adjacency semantic alignment process. In contrast, when $\hat{\gamma}$ is introduced, the F1 score shows a gradual upward trend throughout the iterations. This demonstrates that the secure adjacency factor $\hat{\gamma}$ helps to filter out unreliable risky sets, preventing the propagation of risky information, which is critical for reliable adjacency alignment. To acquire satisfactory performance, the hyperparameters are set to $k=30$ and ${\lambda}_{\textit{se}}=-1$ by default, as indicated by the results in Fig. 12.
@@ V-F1 Discussions on Hyperparameter of SHFA
Fig. 15 illustrates the effect of the number of structure anchors $n_{a}$ (defined in Section IV-B), with ${\lambda}_{\textit{SHFA}}=0.5$. Each structure anchor corresponds to one domain discriminator. Ideally, $n_{a}$ should match the actual number of typical structural types in the source domain, ensuring that each type corresponds to a unique domain discriminator for optimal hierarchical feature adaptation. Setting $n_{a}$ too small may group features from different structural types together, causing confusion and suboptimal adaptation. Conversely, setting $n_{a}$ too large may assign multiple discriminators to the same structural type, resulting in competition effects and reduced adaptation effectiveness. For both the Aircraft LR→HR and Vehicle LR→HR tasks, the results in Fig 15 show that the F1 and mAP score vary with different $n_{a}$. When $n_{a}$ is set to 0, the SHFA module degenerates into a conventional single-adversarial domain adaptation model, resulting in the lowest F1 and mAP values. As $n_{a}$ increases, introducing more hierarchical discriminators into the SHFA module, both F1 and mAP improve, indicating that assigning features to discriminators based on structural similarity enhances feature adaptation performance. The best results are observed when $n_{a}=5$, suggesting that the number of domain discriminators matches the typical structural types in the source domain, achieving optimal hierarchical feature adaptation. Further increasing $n_{a}$ introduces excessive domain discriminators, leading to competition effects and a slight decline in performance. Therefore, it is recommended to set $n_{a}$ to around 5 to ensure optimal performance.